5. The cost of new common stock

True or False: The following statement accurately describes how firms make decisions related to issuing new common stock.

The cost of issuing new common stock is calculated the same way as the cost of raising equity capital from retained earnings.

False: Flotation costs need to be taken into account when calculating the cost of issuing new common stock, but they do not need to be taken into account when raising capital from retained earnings.

True: The cost of retained earnings and the cost of new common stock are calculated in the same manner, except that the cost of retained earnings is based on the firm’s existing common equity, while the cost of new common stock is based on the value of the firm’s share price net of its flotation cost.

Alpha Moose Transporters is considering investing in a one-year project that requires an initial investment of $500,000. To do so, it will have to issue new common stock and will incur a flotation cost of 2.00%. At the end of the year, the project is expected to produce a cash inflow of $550,000. The rate of return that Alpha Moose expects to earn on its project (net of its flotation costs) is    (rounded to two decimal places).

Sunny Day Manufacturing Company has a current stock price of $33.35 per share, and is expected to pay a per-share dividend of $2.45 at the end of the year. The company’s earnings’ and dividends’ growth rate are expected to grow at the constant rate of 5.20% into the foreseeable future. If Sunny Day expects to incur flotation costs of 5.00% of the value of its newly-raised equity funds, then the flotation-adjusted (net) cost of its new common stock (rounded to two decimal places) should be   .

Alpha Moose Transporters Co.’s addition to earnings for this year is expected to be $420,000. Its target capital structure consists of 35% debt, 5% preferred, and 60% equity. Determine Alpha Moose Transporters’s retained earnings breakpoint:

$665,000

$700,000

$805,000

$840,000
Answer #1

1) Initial investment = $500,000
Floatation Costs = 2%
Expected cash inlfow after a year = $550,000

Initial Investment (Net of floatation costs) =Initial investment * (1 - Floatation Costs)
= 500,000 * (1 - 0.02)
= 500,000 * 0.98
Initial Investment (Net of floatation costs) = $490,000

Returns required = [Expected cash inflow after a year / Initial Investment (Net of floatation costs)] - 1
= (550000 / 490000) - 1
Returns required = 0.12244897 or 12.2449%

2) Cost of new equity (net of floatation costs) = {D1 / [P0 * (1 - F)]} + G
where D1 = Dividend to be paid in next period
P0 = Price today
F = Floatation costs
G = Growth Rate

Cost of new equity (net of floatation costs) = {2.45 / [33.35 * (1 - 5%)]} + 5.2%
= [2.45 / (33.35 * 0.95)] + 0.052
= 2.45 / 31.6825 + 0.052
= 0.0773297561745 + 0.052
Cost of new equity (net of floatation costs) = 0.1293297561745 or 12.9329%

3) Retained Earnings = 420,000
Weightage of Equity = 60%

Retained Earnings breakpoint = Retained Earnings / Weightage of Equity in Capital Structure
= 420,000 / 0.6
Retained Earnings breakpoint = $700,000

Hence, Option B is the right answer.